Attending: Eric, Iwona, Chris Powell, Clayton, Mike, Lisa
Account Request Form: Discussed the new form that Clayton has been working on. It features things like drop-down menus for Institution and Country and also automates data entry into NIM so all the PI has to do is approve or disapprove the account request. Also discussed who gets email when the request is made. Currently it goes to Eric and a contact for each group. The new form sends email to Eric and all PI's, Pi Proxies, and Project managers for the group. Users didn't express a strong preference for either method so the new form will work in this manner initially at least.
Cluster Status/Utilization: Utilization has often not been at full capacity recently for a number of reasons including Alice not running much, STAR not having eliza17 and rolling upgrades that require draining nodes.
Outages: eliza17 is still down. There has been a lot of work with Dell and the hope is that it will recover in the next couple days. /project was also down for a day or so last week.\
Upcoming Downtimes: There is some NERSC downtime on May 8th but PDSF will remain up.
Grid: ATLAS issues with expired CRL's have been sorted out. The grid server swap is still in progress. BDII is getting installed on pdsfgrid1.
SL6 testing: Lisa sees no mysql in 6.2 - suggestion is to use 6.2 instead.